# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v1

@@ 5.2.1 Baselines for comparison
  • Data Escrow: We implemented a private query escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use Algorithm 1 to find $\epsilon$ automatically. We simulated the protocol where the data controller registers their dataset to the escrow, creates a policy that allows the data analyst to issue queries to their data using the private query application, and the analyst issues their query and receives the query output.
@@
  • Standalone without DP: This baseline simply runs the given query without DP. The dataset is in plaintext.
  • Standalone with DP: This baseline represents the central DP model. It runs the given query and applies the Laplace Mechanism to the query output. The dataset is in plaintext.
@@
5.2.2 Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD’s Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
5.2.3 Dataset
@@ 5.2.4 Queries
  • Q1: Number of people between age 20 and 30 (simple predicate with range)
  • Q2: Number of people by each race (aggregate & group-by)
  • Q3: Number of people who are female and have income greater than 50K (two simple predicates)
5.2.5 Results
@@
Figure 8: Average query time over 10 runs
@@ 5.3 Microbenchmarks
5.3.1 Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm as a registered application in the escrow.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for size datasets of size 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they are not the same copy of each other. We included all attributes and all records in the dataset.
@@ 5.3.1 Is Find-$\epsilon$-from-PRI practical?
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num=13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==’Asian-Pac-Islander’ AND age>=30 AND age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass==’Federal-gov’ OR workclass==’Local-gov’ or workclass==’State-gov’
@@
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,0.099,\cdots,0.001]$ (37 values in total). We set $\epsilon_{c}=0$, $p=50$ and $\tau=0.01$. We used six parallel processes for computing the PRIs.
Figure 9: Average runtime to find $\epsilon$ over 10 runs
@@
5.3.2 Effect of parameter $p$
To find a suitable $\epsilon$, the algorithm computes the ratio of the maximum PRI over the PRI at the $p$th percentile; if the ratio is close to 1, the PRIs are considered to reflect adequate privacy protection by the controller’s privacy requirement. Specifically, we are interested in choosing the 00th (minimum), $25$th, and $50$th (medium) percentile. The smaller $p$ is, the stricter the privacy requirement and the harder it is to find $\epsilon$, meaning more noise needs to be added to achieve a ratio closer to 1.
@@ 5.3.2 Effect of parameter $p$
Setup. We used the Adult dataset with 100000 rows that we generated in Section 5.3.1, and we used the same settings (except $p$). We ran the five queries in Section 5.3.1 10 times and we report the $\epsilon$ selected by our algorithm in each run.
Figure 10: $\epsilon$ found when using the PRI at the 0, 25, and 50th percentile to compare against the maximum PRI
Results. Figure 10 shows the $\epsilon$ found when $p=0,25,50$. As expected, using higher percentiles tends to find larger $\epsilon$ than lower percentiles, because less noise needs to be added to bring the maximum PRI closer to the PRI at the larger percentile. When using the minimum PRI, Q1 – Q4 find very small $\epsilon$ that are between 0 and 1. Q1 and Q3, both counting queries, are able to find larger $\epsilon$ since the queries have a global sensitivity of 1, i.e., the maximum difference between PRIs will also be fixed at 1, it becomes easier to make the ratio closer to 1. Whereas Q2 and Q4, a group-by query and an average query with higher sensitivities, require more noise to be added to bring the PRIs closer. For Q5, using the minimum PRI only finds $\epsilon$ once, since the sensitivity is too large, primarily due to the domain size of fnlwgt), the DP mechanism will not be able to add enough noise of such scale beyond certain $\epsilon$ (when $\epsilon=0.1$, the noise scale will be larger than 10,000,000, which is the threshold we use to determine whether the DP mechanism can be applied). However, when the requirement is relaxed, it is able to find larger $\epsilon$ such that the noise scale does not exceed the threshold.
6 Related Work
@@
Choosing privacy loss budget. Different approaches [30] [26] [23] [36] [39] [40] [28] [48] [32] [28] aim to help users choose $\epsilon$ when using DP. One line of work [39] [40] [28] [48] [32] implements an interface to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$; they use the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of two different parties: the data analyst and the prospective participant who will choose to participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the "cost" of whether to participate in the study, which is impractical for non-experts. In addition, Cummings et al. and Xiong et al. investigate user’s understanding of DP guarantees through multiple user studies, which offer insights on facilitating communicating the privacy implications of DP to users and help them choose $\epsilon$. Jayaraman and Evans evaluates the impact of $\epsilon$ on ML tasks and specifically their privacy-utility tradeoffs. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
@@
Interpreting utility. Tao et al. proposes DPXPlain to explain aggregate query answers under DP by comparing answers with and without DP, and assessing whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers. Moreover, feasibility studies [8][14] have been conducted to evaluate real-world usage of DP.
@@
Table 1: Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. Alternatives beyond the central model aim to deploy DP with fewer trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted data curator and it has been adopted in various real-world applications for collecting client data from their software and devices [24] [18] [17]. In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the data curator; the data curator does not need to be trusted since it only sees the differentially private version of data. However, the primary disadvantage of the local DP model is that it introduces a much larger error than the central DP model. The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees [34] [6], thus it requires less noise to be added when users perturb their data and offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator [20] [42] [9] [4] [45] [46] [49] [10] [44] [11] [12]. We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
7 Conclusions
To ease the deployment of DP, we address challenges faced by controllers and analysts by 1) defining the privacy risk indicator (PRI) that indicates the relative privacy loss of each individual for a given $\epsilon$ and output and an algorithm to choose $\epsilon$ using the PRIs. 2) introducing an utility signaling protocol that helps analysts interpret the impact of DP on their downstream tasks. We propose a platform implementation on top of a data escrow. We demonstrated our contributions through an IRB-approved user study and empirical evaluation with other DP platforms.

### v2

@@ 6.2.1. Baselines for comparison
  • Data Escrow: We implemented an escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use the Find-$\epsilon$-from-PRI algorithm to select a suitable $\epsilon$. We simulate the two entities (controller and analyst) in the escrow. A policy is first created by the controller to allow the query output to be released to the analyst.
@@
  • Standalone with DP: This baseline runs the given query and applies the Laplace Mechanism to the query output in a central platform. The dataset is in plaintext. This represents the central DP deployment scenario without the overhead of additional security measures.
  • Standalone without DP: This baseline simply runs the given query without DP in a central platform. The dataset is in plaintext. We include this baseline to show the overhead of using DP.
@@
6.2.2. Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
6.2.3. Dataset
@@ 6.2.4. Queries
  • Q1: Number of people between age 20 and 30
  • Q2: Number of people by each race
  • Q3: Number of females who have income greater than 50K
6.2.5. Results
@@
Figure 8. Average query time (in log scale) over 10 runs
@@ 6.3. Microbenchmarks
6.3.1. Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm in the data escrow to further demonstrate its performance for computation-heavy tasks.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for datasets of sizes 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all attributes in the dataset.
@@ 6.3.1. Is Find-$\epsilon$-from-PRI practical?
  • Q1: SELECT COUNT(*) FROM adult WHERE income==’>50K’ AND education_num==13 AND age==25
  • Q2: SELECT marital_status, COUNT(*) FROM adult WHERE race==
    ’Asian-Pac-Islander’ AND 30<=age<=40 GROUP BY marital_status
  • Q3: SELECT COUNT(*) FROM adult WHERE native_country!=
    ’United-States’ AND sex==’Female’
  • Q4: SELECT AVG(hours_per_week) FROM adult WHERE workclass in (’Federal-gov’, ’Local-gov’, ’State-gov’)
@@
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total). We set $p=50$. We used six parallel processes for computing the PRIs.
Figure 9. Average runtime (in log scale) to find $\epsilon$ over 10 runs
@@
6.3.2. Effect of parameter $p$
The Find-$\epsilon$-from-PRI algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$. Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer. We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
@@ 6.3.2. Effect of parameter $p$
Setup. We use the Adult dataset with 100000 rows that we generated in Section 6.3.1. We ran each of the five queries from Section 6.3.1 for 50 times, and report the $\epsilon$ chosen by our algorithm in each run.
Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$
Results. Figure 10 shows boxplots of $\epsilon$ found by the algorithm for different $p$. As shown in the figure, setting larger $p$ tends to find larger $\epsilon$. Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 for Q1–4 and less than 1 for Q5. Queries with larger sensitivity (Q2, Q4, Q5) can find larger $\epsilon$ when $p>50$, and queries with sensitivity of 1 (Q1, Q3) tend to find smaller $\epsilon$. Overall, the results empirically demonstrate that the algorithm is useful for finding a suitable $\epsilon$ (and corresponding query output) that is consistent with the controller’s privacy preferences.
7. Related Work
Defining PRI. The PRI shares similar intuition as Personalized Differential Privacy (PDP) (Jorgensen et al., 2015) in that different individuals may need different levels of protection since some records are more sensitive than others. In contrast to PDP, which requires each individual to specify their own privacy requirement through a privacy budget, the PRI is computed by controllers to indicate the realized privacy loss of each individual if a DP output were to be released.
@@ 7. Related Work
Choosing privacy loss budget. Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP. One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ligett et al. empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts. Cummings et al. and Xiong et al. investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
@@
Interpreting utility. Tao et al. proposed DPXPlain to explain aggregate query answers by comparing answers with and without DP and assess whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers.
@@
Table 1. Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. The threat model of the central DP model assumes adversaries cannot breach the platform and force unintended data flows, which is unrealistic. Alternatives beyond the central model aim to deploy DP with different trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted platform and it has been adopted in various real-world applications for collecting client data from their software and devices  (Erlingsson et al., 2014; Ding et al., 2017; Differential Privacy Team, 2017). In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the platform; adversaries can control the platform but they only gain access to DP data. However, LDP also introduces a much larger error ($O(\sqrt{n})$, where n is the number of individuals) than the central DP model ($O(1)$). The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees  (Kasiviswanathan et al., 2011; Balle et al., 2018), thus it requires adding less noise when users perturb their data and it offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator  (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017; Bogatov et al., 2021). We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
8. Conclusions
In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain additional understanding of the effects of applying DP to the results of an analyst’s query. This translates into less burden choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
In addition, we show that it is possible to augment analysts’ understanding of the effects of DP on their results by using a utility signaling protocol that reveals the isolated impact of DP on a downstream task. We then gather the requirements gleaned from the previous two contributions and propose a platform, based on a data escrow, to deploy both ideas.